OpenStudy (anonymous):

An experiment consists of rolling a fair die three times. What is the probability of getting an even number on all three rolls?3/6-1/2

Parth (parthkohli):

P(Getting an even number) = 1/2 Well, just multiply all three if we want to do calculations like these. \(\Large \color{MidnightBlue}{\Rightarrow ({1 \over 3})^3 }\)

Parth (parthkohli):

* \(\Large \color{MidnightBlue}{\Rightarrow ({1 \over 2})^3 }\)

OpenStudy (lgbasallote):

probability of even number in one roll of die is 3/6 = 1/2 so \[\large (\frac{1}{2} )^3\]

OpenStudy (anonymous):

thanks